OpenStudy (anonymous):

Find f^-1 for f(x)=5x-7

OpenStudy (anonymous):

@hoblos

OpenStudy (hoblos):

y=5x-7 to find the inverse you have to solve for x 5x-7=y 5x = y+7 x=(y+7)/5 then switch x and y to get y=(x+7)/5 thus f-1 = (x+7)/5